OpenStudy (anonymous):

The future value of a continuous income stream of $500 per year over the next 10 years, assuming an annual interest rate of 4% compounded continuously is

OpenStudy (anonymous):

\[a)\int\limits_{0}^{20} 500e^(-0.06(20-t))dt\] \[b)\int\limits_{0}^{20} 500e^(-0.06t) dt\] \[c) \int\limits_{0}^{20} 500(1.06)^(20-t)dt\] d) none of the above